OpenStudy (anonymous):

integrate sec^2x/(secx+tanx)^9/2dx

OpenStudy (anonymous):

Is this what you wish to evaluate?\[\huge \int \frac{\sec^2 x}{(\sec x + \tan x)^\frac{9}{2}}dx\]

OpenStudy (anonymous):

Best thing to do is to write it into Wolfram Alfa! http://www.wolframalpha.com/input/?i=interate+sec^2x%2F%28secx%2Btanx%29^%289%2F2%29+dx